OpenStudy (vuriffy):

Use the trapezium rule with three intervals to find an approximation to:

OpenStudy (vuriffy):

@518nad

OpenStudy (vuriffy):

@zepdrix

OpenStudy (518nad):

do u know any languages?

OpenStudy (vuriffy):

Oh, sorry, I didn't put it up.

OpenStudy (vuriffy):

\[\int\limits_{0}^{?12} \sqrt{4+x^2} dx\]

OpenStudy (518nad):

how many intervals

OpenStudy (vuriffy):

Three

OpenStudy (vuriffy):

Intervals should be 0, 4, 8, 12?

OpenStudy (518nad):

those are the points for your intervals

OpenStudy (518nad):

so your intervals are (0,4) (4,8) (8,12)

OpenStudy (vuriffy):

Yes, that's what I mean. Thank you.

OpenStudy (518nad):

just making sure

OpenStudy (518nad):

ok carry on

OpenStudy (vuriffy):

rad(4+0^2) + rad(4+4^2) * 2 = area for the first one?

zepdrix (zepdrix):

Yes, the 2 should be multiplying the entire thing though, 2 * ( rad(4+0^2) + rad(4+4^2) )

OpenStudy (vuriffy):

That's what I meant, sorry.

OpenStudy (vuriffy):

Soo, I got 79.2 for the final answer when rounded to 3 sig figs. Area(first interval) 12.9443, Area(2nd) 25.4367, Area(3rd) 40.8235

zepdrix (zepdrix):

Integrating will give you 77.977. So yes, your calculations are probably correct \c:/

OpenStudy (vuriffy):

Oh, there's another part to the question, find: \[\int\limits_{?}^{?} 4\cos(\frac{ 1 }{ 3 }x + 2)dx\] Is this supposed to be Integration?

OpenStudy (vuriffy):

It has no limits.

zepdrix (zepdrix):

There is this rule (that I made up) ... I like to call it the Linear Coefficient Rule: If \(\large\rm \frac{d}{dx}F(x)=f(x)\), then,\[\large\rm \int\limits F(ax)=\frac1a f(ax)\]

zepdrix (zepdrix):

This only works with linear coefficients. If the coefficient is on say, an x^2, then something else is going on. But as long as the x is linear, first degree, we can simply divide by the coefficient when we integrate.

zepdrix (zepdrix):

Hopefully this makes a little bit of sense if you think back to your chain rule. When you differentiate,\[\large\rm \frac{d}{dx}\sin(8x)\quad=\quad \cos(8x)\cdot 8\]You get an extra 8 `multiplying` because of the inner function. When you integrate, you LOSE that 8. So you actually end up `dividing it out`.

zepdrix (zepdrix):

\[\large\rm \int\limits \cos(8x)dx=\frac18\sin(8x)\]

zepdrix (zepdrix):

Ok I rambled on for a bit there... Do you understand why this helps though? It allows us to avoid using a substitution.

OpenStudy (vuriffy):

It makes sense, yes. Even with your reasoning, it still equals 12sin(1/3x +2)dx, at least that's what I see from what you showed.

zepdrix (zepdrix):

AHH I wrote my integral backwards, woops.\[\large\rm if~\frac{d}{dx}F(x)=f(x)\] \[\large\rm then~\int\limits f(ax)=\frac1a F(ax)\]Not a big deal though >.<

zepdrix (zepdrix):

So you divided a 1/3 out, which is the same as multiplying by 3? Yayyy good job

OpenStudy (vuriffy):

Oh, yeah, don't you have to add + C?

zepdrix (zepdrix):

Yes yes, that would be a good idea :)

OpenStudy (vuriffy):

Alright wonderful! Thank you so much. I have another question, but it's different. Well, I have two, but I think I can do one on my own, but the other one seems to look confusing. I will tag you?

zepdrix (zepdrix):

I gotta go for a lil bit. You can tag and maybe I'll come by if I find time. Danny boy super smart too though :D

OpenStudy (vuriffy):

Alrighty. That's fine, thank you again!